OpenStudy (anonymous):

A car is traveling at a rate of 40 miles per hour. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 2 hours? In your computations, assume that 1 mile is equal to 1.6 kilometers. Do not round your answers. Rate = km over h Distance traveled in 2 hours = km

OpenStudy (anonymous):

hold up...

OpenStudy (anonymous):

okay please help me tho @FaZe_Rainn

OpenStudy (anonymous):

oh course ;)

OpenStudy (anonymous):

convert kilometer from mph

OpenStudy (anonymous):

40 miles = 64.374 kilometers per hour is that correct ? @FaZe_Rainn

OpenStudy (anonymous):

is it @FaZe_Rainn

OpenStudy (anonymous):

if its is than the car would travel 128.748 kilometers in 2 hours @FaZe_Rainn

OpenStudy (anonymous):

40mphx1.6 = 64km/h distance = 128 km

OpenStudy (johnweldon1993):

The above is correct Rate = \(\large \frac{40miles}{1hour}(\frac{1.6km}{1mile}) = \frac{64km}{hour}\) Now we know that they travel 2 hours...so if it is 64km per hour...times 2 would make it 128 km

OpenStudy (anonymous):

thank you guys so much @johnweldon1993 :)

OpenStudy (johnweldon1993):

Absolutely :D